Circular No. 6586

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) A. J. Apponi, T. C. Pesch, and L. M. Ziurys, Steward Observatory,

University of Arizona; and P. Wehinger and S. Wyckoff, Arizona State University, report on radio observations of comet C/1995 O1 obtained with the NRAO 12-m telescope at Kitt Peak on Mar. 10: “We detected the HNC J=3–2 transition at 271.981 GHz with an integrated line-intensity of 3.89 K km s^{−1} relative to the main beam brightness temperature. On Mar. 9, we independently detected HCO^{+} via the J=2–1 and the J=3–2 transitions at 178.375 and 267.557 GHz, with integrated intensities of 3.11 and 4.12 K km s^{−1}, respectively. Also on Mar. 9, we detected the HCN J=2–1 transition at 177.261 GHz with an integrated intensity of 13.4 K km s^{−1}. Observations of the HCN J=3–2 transition at 265.866 GHz suggest an abundance ratio for [HNC]/[HCN] ∼ 0.5.”

GS 1843+00 R. B. Wilson and B. A. Harmon, Marshall Space Flight Center, NASA;

D. M. Scott, M. H. Finger, and C. R. Robinson, Universities Space Research Association; D. Chakrabarty, Massachusetts Institute of Technology; and T. A. Prince, California Institute of Technology, report for the Compton Observatory BATSE Team: “Pulsations from the transient x-ray binary pulsar GS 1843+00 have been detected since Mar. 3 in the energy range 20–50 keV. Using observations through Mar. 9, the mean barycentric pulse frequency at epoch Mar. 6.0 UT was (3.382592 ± 0.000035) × 10^{−2} Hz. The observed spinup rate during this observation is (4.18 ± 0.12) × 10^{−11} Hz s^{−1}. The mean RMS pulsed flux integrated over 20–50 keV was (3.75 ± 0.24) × 10^{−10} erg cm^{−2} s^{−1} on Mar. 8–9. Earth-occultation measurements indicate a steady increase in total intensity from early March to Mar. 8– 9, with a total flux of 5.1 ± 0.5 × 10^{−9} erg cm^{−2} s^{−1}, giving a pulse fraction of ∼ 7 percent. The thermal bremsstrahlung model (without Gaunt factor) yields a temperature of 23 keV. Pulsed and steady emission are also observed on Mar. 10. Updates to source intensity and pulse frequency are posted twice-weekly for this and all other bright sources detected by

BATSE at the web site http://www.batse.msfc.nasa.gov. Observations

are encouraged at all wavelengths, as no optical counterpart is currently known for this source.”

1997 March 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green